OpenStudy (anonymous):

A rectangular garden is bounded on one side by a house and is to be fenced on the other three sides. Fencing material costs $10 per foot for the side parallel to the house and $20 per foot for the other two sides. What are the dimensions of the garden of largest possible area if $2000 is to be spent for fencing material?

OpenStudy (ganpat):

20 * 2 * b + 10 * l = 2000 If area (l*b)= largest, l=b (Square is the largest rectangle) so 40 b + 10 b = 2000 50 b = 2000 b = 40 ft... so largest are = l * b = 40 * 40 = 1600 sq ft..